OpenStudy (anonymous):

Air travel-Rod is a pilot for Crossland Airways. He computes his flight time against a headwind for a trip or 2900 mi at 5 hr. The fight would take 4 hr and 50 minutes if the headwind were half as great. Find the headwind and the planes air speed.

OpenStudy (anonymous):

Let x = Planes speed in mi / min, y = Headwind speed in mi / min x - y = 2900 / 300 = 9.67 x - 0.5 = 2900 / 290 = 10 => 0.5y = 0.33 => y = 0.66 mi / min = 39.6 mi / hr x = 9.67 + 0.66 = 10.33 mi / min = 619.8 mi / hr

OpenStudy (anonymous):

what method did you use to do this...i do not know why my mind will not wrap around this